A hollow plastic sphere is held below the surface of a fresh-water lake by a cord anchored to the bottom of the lake. The sphere has a volume of 0.650 cubic meters and the tension in the cord is 900 N. 

a) Calculate the buoyant force exerted by the water on the sphere. 

b) what is the mass of the sphere? 

c) the cord breaks and the sphere rises to the surface. When the sphere comes to rest, what fraction of its volume will be submerged?
Answer #1
Concepts and reason

The concept required to solve the given problem is buoyant force.

Calculate the buoyant force by first computing mass in terms of density and volume and then taking the product of mass and acceleration due to gravity.

Calculate the mass of the sphere with the help of equilibrium condition of force.

Calculate the fraction of volume of the sphere submerged by taking the ratio of the volume of sphere submerged and volume of sphere and multiplying it by 100.

Fundamentals

The upward force exerted by a fluid on an object when the object is placed in it is called buoyant force.

Archimedes' Principle states that the upward buoyant force that is exerted on a body immersed in a fluid, whether fully or partially submerged, is equal to the weight of the fluid that the body displaces and acts in the upward direction at the center of mass of the displaced fluid.

The equilibrium condition for the force gives,

ΣF=0\Sigma F = 0

Here, FF is force.

(a)

The expression to calculate the buoyant force is,

FB=ρwaterVsphereg{F_B} = {\rho _{water}}{V_{sphere}}g

Substitute 9.8m/s29.8{\rm{ m / }}{{\rm{s}}^2} for gg, 1000kg/m21000{\rm{ kg / }}{{\rm{m}}^2} for ρwater{\rho _{water}} and 0.650m30.650{\rm{ }}{{\rm{m}}^3} for Vsphere{V_{sphere}} in the above equation.

FB=(1000kg/m2)(0.650m3)(9.8m/s2)=6370N\begin{array}{c}\\{F_B} = \left( {1000{\rm{ kg / }}{{\rm{m}}^2}} \right)\left( {0.650{\rm{ }}{{\rm{m}}^3}} \right)\left( {9.8{\rm{ m / }}{{\rm{s}}^2}} \right)\\\\ = 6370{\rm{ N}}\\\end{array}

(b)

The expression to calculate the mass of the sphere is,

m=FBTgm = \frac{{{F_B} - T}}{g}

Substitute 6370N6370{\rm{ N}} for FB{F_B}, 900N900{\rm{ N}} for TT and 9.8m/s29.8{\rm{ m / }}{{\rm{s}}^2} for gg in the above equation.

m=6370N900N9.8m/s2=558kg\begin{array}{c}\\m = \frac{{6370{\rm{ N}} - 900{\rm{ N}}}}{{9.8{\rm{ m / }}{{\rm{s}}^2}}}\\\\ = 558{\rm{ kg}}\\\end{array}

(c)

When the cord breaks the only forces acting on the sphere will be buoyant force and the weight of the sphere.

The density of the sphere will be,

ρsphere=MVsphere{\rho _{sphere}} = \frac{M}{{{V_{sphere}}}}

Here, MM is the mass of sphere and Vsphere{V_{sphere}} is the volume of sphere.

Substitute 558kg558{\rm{ kg}} for MM and 0.650m30.650{\rm{ }}{{\rm{m}}^3} for VV in the above equation.

ρsphere=558kg0.650m3=858kg/m3\begin{array}{c}\\{\rho _{sphere}} = \frac{{558{\rm{ kg}}}}{{0.650{\rm{ }}{{\rm{m}}^3}}}\\\\ = 858{\rm{ kg / }}{{\rm{m}}^3}\\\end{array}

The volume of the sphere submerged will be,

Vsubmerged=Vsphere(ρsphereρwater){V_{submerged}} = {V_{sphere}}\left( {\frac{{{\rho _{sphere}}}}{{{\rho _{water}}}}} \right)

Substitute 0.650m30.650{\rm{ }}{{\rm{m}}^3} for Vsphere{V_{sphere}}, 858kg/m3858{\rm{ kg / }}{{\rm{m}}^3} for ρsphere{\rho _{sphere}} and 1000kg/m31000{\rm{ kg / }}{{\rm{m}}^3} for ρwater{\rho _{water}} in the above equation.

Vsubmerged=(0.650m3)(858kg/m31000kg/m3)=0.558m3\begin{array}{c}\\{V_{submerged}} = \left( {0.650{\rm{ }}{{\rm{m}}^3}} \right)\left( {\frac{{858{\rm{ kg / }}{{\rm{m}}^3}}}{{1000{\rm{ kg / }}{{\rm{m}}^3}}}} \right)\\\\ = 0.558{\rm{ }}{{\rm{m}}^3}\\\end{array}

The fraction of volume submerged will be,

%=VsubmergedVsphere×100\% = \frac{{{V_{submerged}}}}{{{V_{sphere}}}} \times 100

Substitute 0.558kg/m30.558{\rm{ kg / }}{{\rm{m}}^3} for Vsubmerged{V_{submerged}} and 0.650kg/m30.650{\rm{ kg / }}{{\rm{m}}^3} for Vsphere{V_{sphere}} in the above equation.

%=0.558kg/m30.650kg/m3×100=85.8%\begin{array}{c}\\\% = \frac{{0.558{\rm{ kg / }}{{\rm{m}}^3}}}{{0.650{\rm{ kg / }}{{\rm{m}}^3}}} \times 100\\\\ = 85.8\% \\\end{array}

Ans: Part a

The buoyant force exerted by water on the sphere is 6370N6370{\rm{ N}}.

Part b

The mass of the sphere is 558kg558{\rm{ kg}}.

Part b

The fraction of volume submerged will be 85.8%85.8\% .